# J-PLUS: Reconstructing the Milky Way Disc’s star formation history with twelve-filter photometry

### v1

@@ Abstract
Context. Wide-field, multi-filter photometric surveys provide a competitive dataset for reconstructing the Milky Way’s star formation history (SFH) on Galactic scales and offer complementary perspectives on the assembly of the Galactic disc. The twelve-filter photometric system of the Javalambre Photometric Local Universe Survey (J-PLUS) is well suited for this purpose, as its colours are sensitive to stellar chemical abundances, helping to mitigate the age–metallicity degeneracy inherent in colour–magnitude diagram fitting.
Aims. We aim to recover the SFH of the Milky Way disc, disentangling its chemically distinct components, by combining J-PLUS DR3 photometry with Gaia astrometry. We also assess the potential of isochrone fitting to estimate ages and metallicities for individual stars, using them as proxies for detecting evolutionary trends in the disc populations.
Methods. We fit magnitudes and parallaxes of stars using a Bayesian multiple-isochrone fitting technique. The bright region of the colour–absolute-magnitude diagram ( mag) constrains stellar ages, while the faint region provides an empirical metallicity prior that mitigates the age–metallicity degeneracy. We use PARSEC and BaSTI isochrones, adopting both solar-scaled and -enhanced models.
Results. The recovered SFH reveals two chemically distinct sequences: an -enhanced population forming rapidly between and Gyr ago, enriching from to dex; and a solar-scaled sequence emerging Gyr ago, dominating after Gyr, with slower enrichment that plateaus near solar metallicity by Gyr. Our analysis of individual stars shows that metal-rich () stars are confined to kpc, whereas metal-poor () stars reach kpc.
Conclusions. The simultaneous fitting of solar-scaled and -enhanced isochrones reveals distinct formation epochs for chemically different disc populations. These findings demonstrate the capability of J-PLUS multi-filter photometry, in synergy with Gaia parallaxes, to mitigate age–metallicity degeneracies and to disentangle patterns in the SFHs associated with the thin and thick discs. The methodology is readily applicable to forthcoming wide-field, multi-filter surveys, enabling detailed mapping of the Milky Way’s temporal and chemical evolution.
@@ 7 Data availability
The value-added catalogue containing stellar ages, metallicities, and photometric data used in this work is available at Zenodo (https://zenodo.org/records/17534761DOI: 10.5281/zenodo.17534761)
@@ A.1 APOGEE sample
To ensure a high-quality and reliable sample from the APOGEE DR17 catalog, we selected stars with to guarantee precise stellar parameter and abundance determinations. We required stars to have . Additionally, we required STARFLAG, retained only those stars for which FE_H_FLAG and mas.

### v2

@@ Abstract
Context. Wide-field, multi-filter photometric surveys enable the reconstruction of the Milky Way’s star formation history (SFH) on Galactic scales and provide complementary insights into disc assembly. The twelve-filter system of the Javalambre Photometric Local Universe Survey (J-PLUS) is particularly suitable, as its colours trace stellar chemical abundances and help mitigate the age–metallicity degeneracy in colour–magnitude diagram fitting.
Aims. We aim to recover the SFH of the Milky Way disc and separate its chemically distinct components by combining J-PLUS DR3 photometry with Gaia astrometry. We also test the potential of isochrone fitting to estimate ages and metallicities for individual stars as proxies for disc evolutionary trends.
Methods. We fit magnitudes and parallaxes of stars using a Bayesian multiple-isochrone technique. The bright region of the colour–absolute-magnitude diagram ( mag) constrains stellar ages, while the faint region provides an empirical metallicity prior that mitigates the age–metallicity degeneracy. Both PARSEC and BaSTI isochrones, in solar-scaled and -enhanced versions, are adopted.
Results. The recovered SFH shows two sequences: an -enhanced population forming rapidly between and Gyr ago, enriching from to dex; and a solar-scaled sequence emerging Gyr ago, dominating after Gyr with slower enrichment, reaching solar metallicity by Gyr. Metal-rich () stars are confined to kpc, whereas metal-poor () stars reach kpc.
Conclusions. Simultaneous fitting of solar-scaled and -enhanced isochrones reveals distinct formation epochs for the thin and thick discs. J-PLUS multi-filter photometry, combined with Gaia parallaxes, effectively mitigates age–metallicity degeneracies and enables detailed mapping of the Milky Way’s temporal and chemical evolution.
@@ 7 Data availability
The files containing the inferred values used in Figs. 6, 7, and 8, together with the catalogue containing the individual stellar ages and metallicities used in Figs. 9, 10 and 13, are available at Zenodo (DOI: 10.5281/zenodo.17592911).
@@ A.1 APOGEE sample
To ensure a high-quality and reliable sample from the APOGEE DR17 catalog, we selected stars with to guarantee precise stellar parameter and abundance determinations. We required stars to have Simple. Additionally, we required STARFLAG, retained only those stars for which FE_H_FLAG and mas.